# Ranked from Within: Ranking Large Multimodal Models for Visual Question Answering Without Labels

### v1

@@ 4.3 Uncertainty by Self-Consistency
Another approach involves examining the non-deterministic generations produced by models. The core intuition is that a more accurate model will produce predictions closely aligned with the original answer, while a less accurate model may yield more varied responses with each inference. In the case of LMMs, the temperature parameter controls the randomness of predictions: a temperature of zero forces deterministic predictions, where the model selects only the token with the highest probability. When is larger than , the model generates tokens more stochastically, selecting tokens based on probabilities above a threshold. As increases, the model becomes increasingly stochastic in its token generation [7, 9, 23].
@@
To analyze the consistency in these stochastic predictions, we explore two common methods: BLEU [59] and BERTScore [82]. BLEU is a n-gram-based metric that evaluates the similarity of generated sequences to the reference answer, while BERTScore uses a pre-trained language model to embed answers and measures similarity in embedding space. Then, we use the mean value of similarities to represent the consistency for the input sample, which can be denoted as , where is the number of stochastic inferences, is the similarity function (e.g., BLEU), and and are the -th stochastic prediction and the original answer, respectively. Following the practice outlined in [7, 9, 23], we collect five stochastic inferences per sample and set to maintain a relatively high degree of stochasticity in LMM generation while keeping compute overhead manageable. We adopt a unigram BLEU and denote the methods using BLEU and BERTScore as SampleBLEU and SampleBERT, respectively. Furthermore, BERTScore tends to assign high similarity scores between single letters, such as “A” and “B”, which should be considered very different in the contexts of MCVQ. For example, BERTScore gives a high similarity score of for “A” and “B”, making it ineffective for distinguishing model performance. To address this, we modified the response to include the full answer text (e.g., “A. North America”), denoted as Sample*BERT.
4.4 Uncertainty With a Labeled Proxy Dataset
@@
Beyond the uncertainty score calculated directly on the target dataset, we also explore scores obtained using a proxy labeled dataset. In [18], average thresholded confidence (ATC) has been proposed. ATC calculates a threshold on a validation set (e.g., CIFAR-10 [37] validation set) and considers an image correctly classified on a new dataset with the same task if its confidence score exceeds the threshold. With the derived threshold, model performance on the target domain is estimated by the proportion of samples with confidence scores higher than the threshold. For LMMs, we use an existing benchmark as a proxy set to calculate . The calculation of ATC is formulated as follows:
@@
where is a binary indicator function, represents an uncertainty estimation method, which we use NLLmin here. ATC provides both a ranking of model performance and an estimate of the expected performance.
Method | MCVQ | VQA | ||||||||||||||||
SQA-I | AI2D | MMMU | RWQA | ChartQA | OCRVQA | TextVQA | DocVQA | Average | ||||||||||
AoL | ||||||||||||||||||
NLLF | ||||||||||||||||||
NLLP | ||||||||||||||||||
NLLmin | ||||||||||||||||||
NLLavg | ||||||||||||||||||
EntF | ||||||||||||||||||
EntP | ||||||||||||||||||
Entmax | ||||||||||||||||||
Entavg | ||||||||||||||||||
SampleBLEU | ||||||||||||||||||
SampleBERT | ||||||||||||||||||
Sample*BERT | ||||||||||||||||||
ATC |
@@ 5.1 Experiment setup
We consider both multiple choice visual question (MCVQ) and visual question answering (VQA) benchmarks. MCQ and VQA are both types of question-answering formats for evaluating LMMs. The difference lies on that the former is provided with several answer options, out of which only one (or sometimes more) is correct, the model is only required to select an answer option(s). In contrast, the latter is usually open-ended, models may generate answers freely. We consider widely-adopted MCVQ and VQA benchmarks. They are: (1) the subset of ScienceQA [47] with images (SQA-I) and AI2D [22] which assess LMMs’ scientific knowledge; (2) ChartQA [52], OCRVQA [55], TextVQA [67] and DocVQA [53] to examine their ability to recognize optical character; (3) RealWorldQA (RWQA) [77] which evaluates LMMs’ vision-centric capability; (4) MMMU [80] which assays LMMs on multi-discipline tasks that demand college-level subject knowledge. Note that, SQA-I, AI2D, RWQA and MMMU are MCVQ datasets, while the others are VQA.
@@
We consider two different settings. For the first setting, the goal is to choose the best LMM over all different series of LMMs. We include LLaVA-V1.5 [44], ShareGPT4V [6], LLaVA-NeXT [45], InstructBLIP [11], LLaVA-NeXT-Interleave [42], LLaVA-OneVision [40], Eagle [65], mPLUG-Owl [41], InternVL [8], PaliGemma [5], Mantis [27]. In total, we have collected different models, which all can be accessed on Huggingface [76].
5.2 Main Observations
@@
Accuracy-on-the-line could be a less reliable indicator for ranking LMMs in new domains. Table 1 summarizes the ranking performance of all methods across eight tasks. The AoL performance is calculated as the average correlation strength when using the other seven domains to predict model rankings in the target domain. We observe that AoL does not achieve consistently high correlation with model performance in seven out of the eight tasks. Although AoL shows strong results on RealWorldQA, where it performs best, uncertainty-based methods also demonstrate high correlation strength. For instance, NLLmin exhibits a Spearman’s lower but a weighted Kendall’s higher than AoL. These findings suggest that relying on existing benchmarks alone to select models for safe deployment can be risky, as they may not capture the statistical characteristics of target domains.
The choice of token matters for output-probability-based methods’ ranking capabilities.
@@
In generated sequences, four ways of token usages can be used to extract model uncertainty. For the two on specific positions, i.e., the first and penultimate tokens, their predictive performance depends on the task type (i.e., MCVQ vs. VQA). The uncertainty associated with generating the first token is more indicative for MCVQ tasks, while the penultimate token generally proves more predictive for VQA tasks. This difference is due to the nature of MCVQ tasks, which often prompt models to generate only a single option letter (e.g., “A”), making the first token crucial for evaluating response accuracy. In contrast, VQA tasks require open-form responses consisting of multiple tokens, making the penultimate token-reflecting the model’s understanding of both the question and the answer-more informative.
@@
For the two span within the generated output, i.e., NLLmin and NLLavg, we find they are more stable and less task-specific compared to individual tokens. While NLLmin and Entmax (reflecting the least confident token) is more effective for MCVQ tasks, both methods perform similarly on VQA tasks. The superior performance of using the first token and the least confident token as ranking indicators for MCVQ suggests that a single token (typically the option letter) holds significant meaning. Considering all potential tokens can sometimes result in an unexpectedly higher confidence level, as LMMs may generate the complete answer alongside the option letter (e.g., “B. Columbia”). The tokens following the option letter often have high confidence levels, leading to an overall increase in estimated uncertainty. This tendency results in models generating entire answers with lower uncertainty scores, yielding higher ranks. These findings underscore the importance of further exploration into the optimal token positions for uncertainty estimation in LMMs. One potential approach is leveraging a language model to identify the position of the option letter and use its softmax probability as the uncertainty measure for the complete response.
@@
Negative log likelihood (NLL) is more stable and predictive for model ranking than normalized entropy.
@@
For seven out of eight tasks, NLL consistently shows stronger correlation with model performance than entropy. Both NLL and entropy leverage softmax probabilities during token generation; however, NLL focuses on the maximal probability, while entropy considers all entries. The lower correlation strength of normalized entropy may stem from the significantly different vocabulary sizes of various LMMs. For instance, LLaVA-V1.5-7B has a vocabulary of k tokens, while PaliGemma-3B-mix has a much larger vocabulary of k tokens. Despite being normalized by vocabulary size, entropy may be still susceptible to noise than NLL. Future work could include applying dimension reduction techniques or limiting consideration to the top- most probable tokens or a pre-defined set of tokens.
@@
The performance of sample-based methods is influenced by the nature of the task. They yield higher correlation scores on VQA tasks compared to MCVQ tasks. In VQA, models typically produce more varied responses, making BLEU and BERTScore effective for capturing uncertainty. However, MCVQ tasks constrain models to select from a pre-defined set of answers. While non-zero temperature introduces some randomness, the variation between stochastic inferences is limited. Additionally, we also find that Sample*BERT has higher correlation scores than SampleBERT on four MCVQ benchmarks and suggests that developing more advanced algorithms to capture uncertainty from multiple stochastic inferences could be beneficial.
Although sample-based methods generally show lower overall performance compared to other methods, they remain competitive without the need to examine model architecture or intrinsic states. This highlights their potential for ranking LMMs that are accessed through APIs, such as GPT-4V [1]. Additionally, the performance of sample-based methods can be enhanced by tuning the temperature parameter and increasing the number of stochastic inferences, which could yield more reliable results.
ATC indicates model ranking to some extent, but the correlation strength is influenced by the choice of proxy dataset.
We compute ATC performance as the average correlation when using the other seven datasets as proxy datasets. Our analysis reveals the choice of proxy dataset is critical, as the scale of uncertainty calculated on different datasets can vary. This variation can lead to an inaccurately estimated threshold for determining instance correctness. A potential improvement involves adopting uncertainty calibration methods, such as temperature scaling [20], to calibrate model uncertainty onto a consistent scale.
6 Discussion
@@
NLLmin | |||||
NLLavg | |||||
Entmax | |||||
Entavg | |||||
SampleBLEU | |||||
ATC |
@@
So far, all experiments have focused on selecting LMMs from different model series. However, in some scenarios, the objective is to choose a training recipe that yields better-performing LMMs within the same model series. These models may be trained with additional fine-tuning steps, varied data augmentations, different training sources, or different language models (e.g., Vicuna and Mistral [26]) and visual encoders, such as CLIP and DINOv2 [58]. For our analysis, we use the LLaVA series due to its widespread adoption. Specifically, we employ 15 different LLaVA prismatic (LLaVA-pri) models [31]. We report results on four datasets in Table 2: AI2D, MMMU, TextVQA, and ChartQA, with AI2D and TextVQA being used to evaluate the performance of different LLaVA variants in the original paper. The AoL and ATC performance metrics are computed in the same manner as described in Section 5, using the average correlation across other three datasets.
@@
Our findings align with those from ranking different series of LMMs. First, using model performance on a single existing dataset may not accurately reflect model rankings on a different domain, even when the models are trained with a similar pipeline and minor variations. Second, we observe that ranking different LLaVA-pri models in MCVQ presents a significant challenge, as the variance in model performance on MCVQ tasks is lower compared to VQA tasks. For instance, the performance gap on AI2D is only , whereas the gap on TextVQA exceeds . This indicates that ranking methods must capture subtle differences between models. We find that SampleBLEU remains effective, while NLL and Entropy may not capture these nuances accurately. Additionally, the decrease in correlation between uncertainty estimated by softmax probability suggests that although modifications to the training pipeline may have a slight effect on performance, they can lead to significant variations in the confidence levels for generation. This finding underscores the importance of assessing LMMs more comprehensively, beyond accuracy alone.
Analysis on the weak correlation of AoL.
@@
Figure 2(b) illustrates the correlation of model performance across different benchmarks. We observe that while TextVQA, ChartQA, DocVQA, and OCRVQA all aim to assess the capability of LMMs to recognize optical characters, the correlation between them varies significantly. Specifically, model performance on TextVQA, ChartQA, and DocVQA shows strong correlations, whereas performance on OCRVQA consistently exhibits low correlation with the other three datasets. To explore whether the images or texts within these datasets influence the correlation strength of model performance, we utilize CLIP-L-14-336 [61] to extract image and text embeddings. We then use the Fréchet distance (FD) [15] to measure the similarity between datasets based on these features.
@@
Figure 4 presents a correlation study between the FD of dataset pairs and the correlation strength of model performance on those datasets. We observe a strong correlation for FDs computed using text input features, while FDs measured by image features show a weaker correlation. This suggests that text input similarity is likely a more influential factor for model performance correlation than image similarity. Additionally, orange stars are used to label the points representing the lowest FD for each dataset. Moreover, OCRVQA shows a low correlation strength with other datasets (Fig. 2). The closest dataset in prompt feature space is TextVQA, with a FD of 124, which is considerably higher than the lowest FDs observed between other dataset pairs, typically around 70 or lower. This difference sheds light on the weak model performance correlation between OCRVQA and other OCR-focused datasets.
@@
NLLmin |
Ranking LMMs by a labeled subset of target domain.
@@
Table 3 presents the correlation strength between model performance on labeled instances in the target domain and the overall performance across the entire dataset. We observe that a small labeled set can provides good rankings. Yet, while the sampled data are not representative, such way may not reliably capture the model ranking for the full dataset [60] In contrast, the method NLLmin gives consistently high correlation, highlighting the potential of uncertainty-based methods for effective model ranking.
7 Conclusion

### v2

@@ 4.3 Assessing Uncertainty by Self-Consistency
Another approach involves examining the non-deterministic generations produced by models. The core intuition is that a more accurate model will produce predictions closely aligned with the original answer, while a less accurate model may yield more divergent responses with each inference. In the case of LMMs, the temperature parameter controls the randomness of predictions: a temperature of zero forces deterministic predictions, where the model selects only the token with the highest probability. When is larger than , the model generates tokens stochastically, selecting tokens based on probabilities above a threshold. As increases, tokens are sampled from an increasingly uniform distribution (Chen et al., 2023b; Cobbe et al., 2021).
@@
To analyze the consistency in these stochastic predictions, we explore two common methods: BLEU (Papineni et al., 2002) and BERTScore (Zhang et al., 2019). BLEU is a n-gram-based metric that evaluates the similarity of generated sequences to the reference answer, while BERTScore uses a pre-trained language model to embed answers and measures similarity in embedding space. Then, we use the mean value of similarities to represent the consistency for the input sample, which can be denoted as , where is the number of stochastic inferences, is the similarity function (e.g., BLEU), and and are the -th stochastic prediction and the original answer, respectively. Following the practice outlined in (Chen et al., 2023b; Cobbe et al., 2021; Huang et al., 2023), we collect five stochastic inferences per sample and set to maintain a relatively high degree of stochasticity in LMM generation while keeping compute overhead manageable. We adopt a unigram BLEU and denote the methods using BLEU and BERTScore as Sample and Sample, respectively. Furthermore, BERTScore tends to assign high similarity scores between single letters, such as “A” and “B”, which should be considered very different in the contexts of MCVQ. For example, BERTScore gives a high similarity score of for “A” and “B”, making it ineffective for distinguishing model performance. To address this, we modified the response to include the full answer text (e.g., “A. North America”), denoted as Sample*.
4.4 Assessing Uncertainty With Labeled Proxy Data
@@
Beyond the uncertainty score calculated directly on the target dataset, we also explore scores obtained using a proxy labeled dataset. Garg et al. (2022) proposed average thresholded confidence (ATC), which calculates a threshold on a validation set (e.g., CIFAR-10 (Krizhevsky & Hinton, 2009) validation set) and considers an image correctly classified on a new dataset with the same task if its confidence score exceeds the threshold. With the derived threshold, model performance on the target domain is estimated by the proportion of samples with confidence scores higher than the threshold. For LMMs, we use an existing benchmark as a proxy set to calculate . The calculation of ATC is
@@
where is a binary indicator function, and represents an uncertainty estimation method, for which we use NLL. ATC provides both a ranking of model performance and an estimate of the expected performance.
| Method | MCVQ | VQA | ||||||||||||||||||
| SQA-I | AI2D | RWQA | MMMU | GQA | ChartQA | OCRVQA | TextVQA | DocVQA | Average | |||||||||||
| AoL | ||||||||||||||||||||
| NLL | ||||||||||||||||||||
| NLL | ||||||||||||||||||||
| NLL | ||||||||||||||||||||
| NLL | ||||||||||||||||||||
| Ent | ||||||||||||||||||||
| Ent | ||||||||||||||||||||
| Ent | ||||||||||||||||||||
| Ent | ||||||||||||||||||||
| Sample | ||||||||||||||||||||
| Sample | ||||||||||||||||||||
| Sample* | ||||||||||||||||||||
| ATC | ||||||||||||||||||||
@@ 5 Experiments
In this section, we first introduce the experimental setup including the evaluated datasets and models we considered. Then, we show the results of ranking different large multimodal models (LMMs).
@@ 5.1 Experiment Setup
We choose the task of visual question answering, which is a common way to evaluate LMMs (Liu et al., 2024a; Dai et al., 2023; Beyer et al., 2024; Lu et al., 2022; Yue et al., 2024). We evaluate on multiple choice visual question (MCVQ) and visual question answering (VQA) benchmarks. MCQ and VQA are both types of question-answering formats for evaluating LMMs. For the former, the model is provided with several answer options, out of which the correct subset is to be selected. In contrast, the latter is usually open-ended and the models may generate answers freely. We consider widely-adopted MCVQ and VQA benchmarks. They are (1) the subset of ScienceQA (Lu et al., 2022) with images (SQA-I) and AI2D (Hiippala et al., 2021) which assess LMMs’ scientific knowledge; (2) ChartQA (Masry et al., 2022), OCRVQA (Mishra et al., 2019), TextVQA (Singh et al., 2019) and DocVQA (Mathew et al., 2021) to examine their ability to recognize optical character; (3) RealWorldQA (RWQA) (x.ai, 2024) and GQA (Ainslie et al., 2023) which evaluate LMMs’ vision-centric capability; (4) MMMU (Yue et al., 2024) which assays LMMs on multi-disciplinary tasks that demand college-level subject knowledge. Note that SQA-I, AI2D, RWQA and MMMU are MCVQ datasets, while the others are VQA.
@@
The goal is to choose the best LMM over all different series of LMMs. We include LLaVA-V1.5 (Liu et al., 2024a), ShareGPT4V (Chen et al., 2023a), LLaVA-NeXT (Liu et al., 2024b), InstructBLIP (Dai et al., 2023), LLaVA-NeXT-Interleave (Li et al., 2024b), LLaVA-OneVision (Li et al., 2024a), Eagle (Shi et al., 2024b), mPLUG-Owl (Li et al., 2022), InternVL (Chen et al., 2024), PaliGemma (Beyer et al., 2024), Mantis (Jiang et al., 2024), DeepSeek-VL2 (Wu et al., 2024) and Qwen2-VL (Wang et al., 2024b). In total, we collect different models, which all can be accessed on Hugging Face (Wolf et al., 2020).
5.2 Key Findings
Accuracy-on-the-Line is unreliable for ranking LMMs in new domains.
@@ 5.2 Key Findings
Table 1 summarizes the ranking capability of all methods across eight benchmarks. The AoL performance is calculated as the average correlation strength when using the other seven domains to predict model rankings in the target domain. We observe that AoL does not achieve consistently high correlation with model performance in seven out of the eight benchmarks. Although AoL shows strong results on RealWorldQA, where it performs best, uncertainty-based methods also demonstrate high correlation strength. For instance, NLL exhibits a Spearman’s lower but a weighted Kendall’s higher than AoL. These findings suggest that relying on existing benchmarks alone to select models for deployment could be risky and unstable, as they may not well capture the statistical characteristics of the new target domain.
The choice of token matters for output probability-based ranking.
@@
We studied four ways of using tokens for estimating model uncertainty. For the two on specific positions, the first and penultimate tokens, their predictive performance depends on the task type (i.e., MCVQ vs. VQA). The uncertainty associated with generating the first token is more indicative for MCVQ tasks, while the penultimate token generally proves more predictive for VQA tasks. This difference is due to the nature of MCVQ tasks, which often prompt models to generate only a single option letter (e.g., “A”), making the first token crucial for evaluating response accuracy. In contrast, VQA tasks require open-form responses consisting of multiple tokens, making the penultimate token—reflecting the model’s understanding of both the question and the answer—more informative.
@@
For the two that consider every token in the generated output, i.e., NLL and NLL, we find that they are more stable and less task-specific compared to variants that consider individual tokens. While NLL and Ent (reflecting the least confident token) is more effective for MCVQ tasks, both methods perform similarly on VQA tasks. The higher ranking correlation of using the first token and the least confident token as ranking indicators for MCVQ suggests that a single token (typically the option letter) holds significant meaning. Considering all potential tokens can sometimes result in an unexpectedly higher confidence level, as LMMs may generate the complete answer alongside the option letter (e.g., “B. Columbia”). The tokens following the option letter often have high confidence levels, leading to an overall increase in estimated uncertainty. This tendency results in models generating entire answers with lower uncertainty scores, yielding higher ranks. These findings underscore the importance of further exploration into the optimal token positions for uncertainty estimation in LMMs. One potential approach is leveraging a language model to identify the position of the option letter and use its softmax probability as the uncertainty measure for the complete response.
@@
The negative log-likelihood (NLL) is more stable and predictive for LMMs ranking than normalized entropy.
@@
For seven out of eight benchmarks, NLL consistently shows stronger correlation with model performance than entropy. Both NLL and entropy make use of the softmax probability distribution predicted for each token during generation. However, the NLL-based approaches only consider the maximum probability in this distribution, while the entropy considers all entries. The lower correlation strength of normalized entropy may therefore stem from the significantly different vocabulary sizes of various LMMs. For instance, LLaVA-V1.5-7B has a vocabulary of k tokens, while PaliGemma-3B-mix has a much larger vocabulary of k tokens. Despite being normalized by vocabulary size, entropy may still be more susceptible to noise than NLL. Future work could include applying dimension reduction techniques or limiting consideration to the top- most probable tokens or a pre-defined set of tokens.
@@
The correlation strength of sample-based methods is influenced by the nature of the task. They yield higher correlation scores on VQA tasks compared to MCVQ tasks. In VQA, models typically produce more varied responses, making BLEU and BERTScore effective for capturing uncertainty. However, MCVQ tasks constrain models to select from a pre-defined set of answers. While non-zero temperature introduces some randomness, the variation between stochastic inferences is limited. Additionally, we also find that Sample* has higher correlation scores than Sample on four MCVQ benchmarks and suggests that developing more advanced algorithms to capture uncertainty from multiple stochastic inferences could be beneficial.
Although sample-based methods show weaker overall correlations, they remain competitive without requiring access to model architecture or internal states. This highlights their potential to rank API-based or closed-source LMMs (e.g., GPT-4V (Achiam et al., 2023)).
ATC can be used for LMMs ranking, but the correlation strength is influenced by the choice of proxy dataset.
We compute ATC performance as the average correlation when using the other seven datasets as proxy datasets. Our analysis reveals the choice of proxy dataset is critical, as the scale of uncertainty calculated on different datasets can vary. This variation can lead to an inaccurately estimated threshold for determining instance correctness. A potential improvement involves adopting uncertainty calibration methods, such as temperature scaling (Guo et al., 2017), to calibrate model uncertainty onto a consistent scale.
6 Analysis
This section includes three distinct analyses. The first specifically investigates whether the considered methods are effective for ranking models within the same series. We use LLaVA models as a case study because they are widely adopted and representative of LMM architectures. The other two analyses consider models from different series, consistent with the broader evaluation in Section 5.
@@ 6 Analysis
| NLL | |||||
| NLL | |||||
| Ent | |||||
| Ent | |||||
| Sample | |||||
| ATC | |||||
@@
So far, all experiments have focused on selecting LMMs from different model series. However, in some scenarios, the objective is to choose a training recipe that yields better-performing LMMs within the same model series. These models may be trained with additional fine-tuning steps, varied data augmentations, different training sources, or different language models (e.g., Vicuna and Mistral (Jiang et al., 2023)) and visual encoders, such as CLIP and DINOv2 (Oquab et al., 2023). For our analysis, we use the LLaVA series due to its widespread adoption. Specifically, we employ different LLaVA prismatic (LLaVA-pri) models (Karamcheti et al., 2024). We report results on four datasets in Table 2: AI2D, MMMU, TextVQA, and ChartQA, with AI2D and TextVQA being used to evaluate the performance of different LLaVA variants in the original paper. The AoL and ATC performance metrics are computed in the same manner as described in Section 5, using the average correlation across other three datasets.
@@
Our findings align with those from ranking different series of LMMs. First, using model performance on a single existing dataset may not accurately reflect LMMs rankings on a different domain, even when the models are trained with a similar pipeline and minor variations. Second, we observe that ranking different LLaVA-pri models in MCVQ presents a significant challenge, as the variance in model performance on MCVQ tasks is lower compared to VQA tasks. For instance, the performance gap on AI2D is only , whereas the gap on TextVQA exceeds . This indicates that ranking methods must capture subtle differences between models. We find that Sample remains effective, while NLL and Entropy may not capture these nuances accurately. Additionally, the decrease in correlation between uncertainty estimated by softmax probability suggests that although modifications to the training pipeline may have a slight effect on performance, they can lead to significant variations in the confidence levels for generation. This finding underscores the importance of assessing LMMs more comprehensively, beyond accuracy alone.
Analysis of the weak correlation of AoL.
@@
Figure 2(b) illustrates the correlation of model performance across different benchmarks. We observe that while TextVQA, ChartQA, DocVQA, and OCRVQA all aim to assess the capability of LMMs to recognize optical characters, the correlation between them varies significantly. Specifically, model performance on TextVQA, ChartQA, and DocVQA shows strong correlations, whereas performance on OCRVQA consistently exhibits low correlation with the other three datasets. To explore whether the images or texts within these datasets influence the correlation strength of model performance, we utilize CLIP-L-14-336 (Radford et al., 2021) to extract image and text embeddings. We then use the Fréchet distance (FD) (Fréchet, 1957) to measure the similarity between datasets based on these features.
@@
Figure 4 presents a correlation study between the FD of dataset pairs and the correlation strength of model performance on those datasets. We observe a strong correlation for FDs computed using text input features, while FDs measured by image features show a weaker correlation. This suggests that text input similarity is likely a more influential factor for model performance correlation than image similarity. Additionally, orange stars are used to label the points representing the lowest FD for each dataset. Moreover, OCRVQA shows a low correlation strength with other datasets (Figure 2). The closest dataset in prompt feature space is TextVQA, with a FD of , which is considerably higher than the lowest FDs observed between other dataset pairs, typically around or lower. This difference sheds light on the weak model performance correlation between OCRVQA and other OCR-focused datasets.
@@
| NLL | ||||
Ranking LMMs by a labeled subset of the target domain.
@@
Table 3 presents the correlation strength between model performance on labeled instances in the target domain and the overall performance across the entire dataset. We observe that a small labeled set can provide a good ranking. However, such an approach may not fully capture the model ranking across the entire dataset, since the sampled data may not be representative (Polo et al., 2024). In contrast, NLL gives a more stable correlation, highlighting the potential of uncertainty-based methods for effective model ranking without data annotation.
7 Conclusion and Discussion